OpenStudy (anonymous):

Lagrange Multiplier: f(x,y) = 5x^2 + 2y^2 xy=3

OpenStudy (anonymous):

image included

OpenStudy (anonymous):

i don't understand why it isn't 399/10 i got x=\[(2\sqrt{3} )/ 5\] y=\[(5\sqrt{3} )/ 2\]

OpenStudy (anonymous):

which multiplied together = 3 so then you would plug x and y back into f(x,y) = 399/10

OpenStudy (anonymous):

it's really x = above, y = above and x = - <(negative)> above, y = - <(negative)> above but either one with come out with the same minimum value

OpenStudy (zarkon):

399/10 =39.9 what if we just use x=1 and y=3 then x*y=3 and \[f(x,y) = 5x^2 + 2y^2\] \[f(1,3) = 5\cdot 1^2 + 2\cdot 3^2=5+18=23\] so you don't have the min

OpenStudy (zarkon):

looks like the min is \(6\sqrt{10}\)

OpenStudy (anonymous):

hrm but you can't just plug numbers in like that right?? because you have to solve the system?

OpenStudy (zarkon):

I'm just showing you that your answer is not correct

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so you can do test cases like that?

OpenStudy (zarkon):

Just by observation I could tell that your answer did not give the min

OpenStudy (zarkon):

why don't you show your work and we will tell you where your mistake is

OpenStudy (anonymous):

yeah just kinda hard to put it into the computer

OpenStudy (turingtest):

\[\nabla f(x,y,z)=\lambda\nabla g(x,y,z)\]I'll try to see ho far I can go before I mess up, I'm not always great with these

OpenStudy (anonymous):

yeah, this stuff isn't really that hard conceptually but i have big time trouble with finding all the solutions to systems of equations

OpenStudy (anonymous):

k so it will come out with 10x=\[\lambda\]y 4x=\[\lambda\]x xy=3

OpenStudy (anonymous):

er

OpenStudy (turingtest):

\[\nabla f=<10x,4y>\]\[\lambda\nabla g=<\lambda x,\lambda y>\]\[10x=\lambda x\implies\lambda=10\]\[4y=\lambda y\implies\lambda=4\]ok I seem to be confused...

OpenStudy (anonymous):

x= 3/y plugging into first: 10(3/y) = z * y z = 30/y^2 plugging both into second: y^3 = 22.5 y = 2.823

OpenStudy (turingtest):

x=0 y=0 ?

OpenStudy (turingtest):

obviously not...

OpenStudy (anonymous):

oh you went for lambda first

OpenStudy (zarkon):

\[10x=\lambda y\] \[10x^2=\lambda xy=3\lambda \] \[4y=\lambda x\] \[4y^2=\lambda xy=3\lambda\] \[\Rightarrow 10x^2=4y^2\]

OpenStudy (zarkon):

you don't need to know lambda

OpenStudy (turingtest):

I know that, but I couldn't eliminate it

OpenStudy (anonymous):

yeah, so just eliminate it first

OpenStudy (turingtest):

I was just demonstrating that I was lost

OpenStudy (turingtest):

oh I messed up the system, duh!

OpenStudy (zarkon):

yep...that would cause a problem ;)

OpenStudy (anonymous):

ah k so sqrt 10 * x = 2 * y

OpenStudy (anonymous):

so then you use that with the last problem so basically just get rid of lambda then use the resulting equations 10x^2 - 2*y = 0 x*y = 3

OpenStudy (zarkon):

\[\sqrt{10}x=\pm2y\]

OpenStudy (zarkon):

\[y=\pm\frac{\sqrt{10}}{2}x\]

OpenStudy (anonymous):

thank you both sooo much

OpenStudy (turingtest):

yeah, I'm not going to restate Zarkon, I'm tired anyway g'night y'all

OpenStudy (anonymous):

night